# Provably Correct Automata Embeddings for Optimal Automata-Conditioned Reinforcement Learning

### v1

@@ 1 Introduction
Formal specifications have been proposed for specifying tasks to goal-conditioned policies. While their well-defined semantics make them appealing, approaches that rely on hierarchical planning, i.e., planning over the induced automaton of a formal specification and instructing a goal-conditioned policy to execute the plan ([Jothimurugan et al.(2021)Jothimurugan, Bansal, Bastani, and Alur, Qiu et al.(2023)Qiu, Mao, and Zhu]), are inherently suboptimal due to the myopia of their high-level planner and goal-conditioned policy. Conditioning reinforcement learning (RL) policies on Linear Temporal Logic (LTL) specifications was proposed by [Vaezipoor et al.(2021)Vaezipoor, Li, Icarte, and Mcilraith], using a graph neural network (GNN) to encode abstract syntax trees of LTL formulas. But, generalization is an inherent limitation due to the use of LTL ([Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]).
@@
In this work, we present a theoretical framework for the DFA-conditioned RL problem and show that it is Probably Approximately Correct (PAC)-learnable. As we showed in ([Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]), learning to encode the DFAs while simultaneously learning a policy is challenging due to the sparse reward specified by DFA acceptance. To this end, in the same work, we demonstrated that pretraining and freezing DFA embeddings before applying them to a downstream learning task greatly improve learning efficiency. Therefore, in order for our PAC learnability guarantee to be useful in practice, one must show that such guarantees hold w.r.t. such pretrained and frozen DFA embeddings. To address this, we present a novel method for learning provably correct DFA embeddings, guaranteeing optimal DFA-conditioned RL. We first observe that bisimilar DFAs represent the same task and then use bisimulation metrics, a relaxation of the notion of a bisimulation relation, to embed unique tasks to unique latent representation. Our experimental evaluation shows that the correctness of the learned DFA embeddings improves downstream policy learning.
@@
PAC-learnability of RL objectives given by formal specifications has been studied before. [Yang et al.(2021)Yang, Littman, and Carbin] prove that the optimal policy for any LTL formula is PAC-learnable if and only if the formula can be checked within a finite horizon. A similar result by [Alur et al.(2022)Alur, Bansal, Bastani, and Jothimurugan] shows that without additional information on the transition probabilities, such as the minimum nonzero transition probability, LTL is not PAC-learnable. Later, a positive result for discounted LTL was given by [Alur et al.(2023)Alur, Bastani, Jothimurugan, Perez, Somenzi, and Trivedi]. Our PAC-learnability result can be considered as a multi-task generalization of these previous results, where the policy must satisfy a class of specifications, not a single objective. See [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia] for a more detailed literature review on using formal specification in RL.
[Ferns et al.(2004)Ferns, Panangaden, and Precup] showed that bisimulation metrics can be computed as unique fixed points of a contraction map. A special case of this result for deterministic MDPs and on-policy samples was proved by [Castro(2020)]. Later, [Zhang et al.(2020)Zhang, McAllister, Calandra, Gal, and Levine] used bisimulation metrics to learn invariant observation embeddings for RL, where they proved these metrics can be computed while learning a policy. We will use these results in Section 4 to learn provably correct DFA embeddings. To our knowledge, no prior work considered using bisimulation metrics for learning task representations.
@@ 2 Background
Given a distribution $\mathcal{X}$, we denote its support by $X\triangleq\operatorname{supp}(\mathcal{X})$. Similarly, given a set $X$, $\mathcal{X}\in\Delta(X)$ denotes some distribution over it, where $\Delta(X)\subset X\to[0,1]$ is the set of all distributions over $X$. $\mathbb{I}\{\}$ is the event indicator function, i.e., $\mathbb{I}\{p\}=1$ if and only if $p$ is true, $\mathbb{I}\{p\}=0$ otherwise.
@@ 2.2 Deterministic Finite Automata
We use deterministic finite automaton (DFA) with three-valued semantics as a task representation.
Definition 2
A deterministic finite automaton (DFA) is defined as the tuple $\mathcal{A}=\langle Q,\Sigma,\delta,q_{0},F\rangle$, where $Q$ is the finite set of states, $\Sigma$ is the finite alphabet, $\delta:Q\times\Sigma\to Q$ is the transition function, where $\delta(q,a)=q^{\prime}$ denotes a transition to a state $q^{\prime}\in Q$ from a state $q\in Q$ by observing a symbol $a\in\Sigma$, $q_{0}\in Q$ is the initial state, and $F=F^{\top}\cup F^{\bot}\subseteq Q$ is the set of accepting states. The semantics of a DFA is defined by its extended (lifted) transition function $\delta^{*}:Q\times\Sigma^{*}\to Q$, where
@@
If $\delta^{*}(q_{0},w)\in F^{\top}$, then we say that $\mathcal{A}$ accepts $w$. Similarly, $\delta^{*}(q_{0},w)\in F^{\bot}$, then we say that $\mathcal{A}$ rejects $w$. $\mathcal{A}$ is called a plan DFA if its final states are sink states, i.e., $\forall q\in F,\forall a\in\Sigma.\delta(q,a)=q$.
Remark 1
DFAs can be minimized to a canonical form up to a state isomorphism using Hopcroft’s algorithm [Hopcroft(1971)], denoted by $\operatorname{minimize}(\mathcal{A})$. Notice that all minimized co-safety DFAs with non-empty languages have a single accepting state, denoted by $\top$, as their accepting states are all sink states. We denote the single state accepting DFA by $\mathcal{A}_{\top}$ and the rejecting one by $\mathcal{A}_{\bot}$.
We use DFAs to represent temporal tasks, which can be understood as a plan, encoding the long-term behavior of the policy. However, one can define multiple DFAs for the same task, i.e., DFAs, without any additional assumptions, are not canonical task representations, addressed next.
@@
In what follows, unless stated otherwise, all DFAs are minimized plan DFAs.
We will also need a notion of similarity between DFAs so that we can compare the tasks represented by them. To this end, we define bisimulation relation over DFAs next.
@@ 3 DFA-Conditioned Reinforcement Learning
We will define the DFA-conditioned RL problem over a given distribution of DFAs, similar to the GCRL problem. However, we need some extra structure over the DFAs in this distribution since in our case, goals are not simply sets of states but are DFAs encoding temporal tasks given to the policy.
@@
$\forall\mathcal{A}=\langle Q,\Sigma,\delta,q_{0},F\rangle\in D_{\Sigma,n}.\forall w\in\Sigma^{*}.\operatorname{minimize}(\mathcal{A}^{\prime}=\langle Q,\Sigma,\delta,\delta^{*}(q_{0},w),F\rangle)\in D_{\Sigma,n}.$
@@
  • $D_{\Sigma,n}$ is the set of DFAs,
  • $\Sigma$ is the shared alphabet,
@@
    $R_{D_{\Sigma,n}}(\mathcal{A}_{t},a)=\begin{cases}1&\text{if }T_{\Sigma,n}(\mathcal{A}_{t},a)=\mathcal{A}_{\top}\\ -1&\text{if }T_{\Sigma,n}(\mathcal{A}_{t},a)=\mathcal{A}_{\bot}\\ 0&\text{otherwise.}\end{cases}$
A DFA space is a set of DFAs closed under random walks, i.e., taking any random path in a DFA from this set and pruning the unreachable states results in a DFA in the set. That is, one cannot get a DFA outside of this set by taking a random walk with minimization, hence the naming space.
@@
We now have all the theoretical machinery needed to formally state the DFA-conditioned RL problem. We first define the environment model and then continue with defining the problem.
@@
$\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}=\langle S\times D_{\Sigma,n},A,L,P_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}},R_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}},\iota_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}},\gamma\rangle$
@@
  • $S\times D_{\Sigma,n}$ is the state space,
  • $A$ is the action space,
  • $L:S\to\Sigma$ is the labeling function,
  • $P_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}:(S\times D_{\Sigma,n})\times A\to\Delta(S\times D_{\Sigma,n})$ is the transition probability function defined by
    $P_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}\left((s,\mathcal{A}),a,(s^{\prime},\mathcal{A}^{\prime})\right)=P(s,a,s^{\prime})\mathbb{I}\left\{\mathcal{A}^{\prime}=T_{D_{\Sigma,n}}\left(\mathcal{A},L(s^{\prime})\right)\right\},$
  • $R_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}:S\times D_{\Sigma,n}\times A\to\{-1,0,1\}$ is the reward function defined by
    $R_{D_{\Sigma,n}}(s,\mathcal{A},a)=\begin{cases}1&\text{if }T_{D_{\Sigma,n}}(\mathcal{A},L(s^{\prime}))=\mathcal{A}_{\top}\\ -1&\text{if }T_{D_{\Sigma,n}}(\mathcal{A},L(s^{\prime}))=\mathcal{A}_{\bot}\\ 0&\text{otherwise,}\end{cases}$
@@
  • $\iota_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}\in\Delta(S\times D_{\Sigma,n})$ is the initial state distribution defined by
    $\iota_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}(s,\mathcal{A})=\iota(s)\mathcal{D}_{\Sigma,n}(\mathcal{A}),$
@@
  • $\gamma\in[0,1)$ is the discount factor.
@@
Given an DFA-conditioned MDP $\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}$ as defined in 5, a DFA-conditioned policy is a mapping
@@
$J_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}(\pi)=\mathbb{E}_{(s_{0},\mathcal{A}_{0})\sim\iota_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}}\left[\sum_{t=0}^{\mathcal{A}_{t}=\mathcal{A}_{\top}\lor\mathcal{A}_{t}=\mathcal{A}_{\bot}}\gamma^{t}\,R_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}(s_{t},\mathcal{A}_{t},a_{t})\right],$
@@
$a_{t}\sim\pi(s_{t},\mathcal{A}_{t}),\quad s_{t+1}\sim P(s_{t},a_{t}),\quad\mathcal{A}_{t+1}=T\left(\mathcal{A}_{t},\,L(s_{t+1})\right),$
until accepting or rejecting DFA is reached, i.e., $\mathcal{A}_{t}=\mathcal{A}_{\top}\lor\mathcal{A}_{t}=\mathcal{A}_{\bot}$. The objective is to solve
$\pi^{*}=\arg\max_{\pi}J_{\mathcal{M}\mid\mathcal{M}_{D_{\Sigma,n}}}(\pi),$
i.e., to learn a policy that maximizes the probability of satisfying the temporal specification (by driving the DFA to $\mathcal{A}_{\top}$) while operating in the given underlying MDP dynamics.
Notice the difference between the GCRL problem formulation given in 11 and the DFA-conditioned one given above. Specifically, in GCRL, goals are static, i.e., they are given at the beginning, and the policy conditions on the same goal until it is accomplished. In the DFA-conditioned RL setting, through the labeling function, a given DFA task also evolves (based on the transition dynamics of its DFA space) as the policy interacts with the underlying MDP. Therefore, the policy is essentially learning to navigate two MDPs: the underlying MDP and the DFA space.
3.1 PAC-Learnability of DFA-Conditioned Reinforcement Learning Problem
We show that the DFA-conditioned RL problem is probably approximately correct (PAC) learnable. To do so, we will use the PAC-MDP framework introduced by [Strehl et al.(2006)Strehl, Li, Wiewiora, Langford, and Littman]. An RL algorithm is called PAC-MDP (PAC in MDPs) if it finds a near-optimal policy with high probability in any MDP after a number of interactions that is polynomial in the problem’s key parameters, stated next.

### v2

@@ 1 Introduction
Formal specifications have been proposed for specifying tasks to goal-conditioned policies. While their well-defined semantics make them appealing, approaches that rely on hierarchical planning, i.e., planning over the induced automaton of a formal specification and instructing a goal-conditioned policy to execute the plan ([Jothimurugan et al.(2021)Jothimurugan, Bansal, Bastani, and Alur, Qiu et al.(2023)Qiu, Mao, and Zhu]), are inherently suboptimal due to the myopia of their goal-conditioned policies. Conditioning Reinforcement Learning (RL) policies on Linear Temporal Logic (LTL) specifications was proposed by [Vaezipoor et al.(2021)Vaezipoor, Li, Icarte, and Mcilraith], using a Graph Neural Network (GNN) to encode abstract syntax trees of LTL formulas. However, generalization is an inherent limitation due to their use of LTL ([Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia]).
@@
In this work, we present a theoretical framework for the DFA-conditioned RL problem and show that it is Probably Approximately Correct (PAC)-learnable. As we showed in [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia], learning to encode the DFAs while simultaneously learning a policy is challenging due to the sparse reward specified by DFA acceptance. To this end, in the same work, we demonstrated that pretraining and freezing DFA embeddings and then passing these embeddings to a downstream policy greatly improve learning efficiency. Therefore, in order for our PAC learnability guarantee to be useful in practice, one must show that such guarantees hold w.r.t. such pretrained and frozen DFA embeddings. To address this, we present a novel method for learning provably correct DFA embeddings, guaranteeing optimal DFA-conditioned RL. We first observe that bisimilar DFAs represent the same task and then use bisimulation metrics, a relaxation of the notion of a bisimulation relation, to embed unique tasks to unique latent representations. Our experimental evaluation shows that the correctness of the learned DFA embeddings improves downstream policy learning.
@@
PAC-learnability of RL objectives given by formal specifications has been studied before. [Yang et al.(2021)Yang, Littman, and Carbin] proved that the optimal policy for any LTL formula is PAC-learnable if and only if the formula can be checked within a finite horizon. A similar result by [Alur et al.(2022)Alur, Bansal, Bastani, and Jothimurugan] shows that without additional information on the transition probabilities, such as the minimum nonzero transition probability, LTL is not PAC-learnable. Later, a positive result for discounted LTL was given by [Alur et al.(2023)Alur, Bastani, Jothimurugan, Perez, Somenzi, and Trivedi]. Our PAC-learnability result can be considered a multi-task generalization of these previous results, where the policy must satisfy a class of specifications, not a single objective. See [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia] for a more detailed literature review on using formal specification in RL.
[Ferns et al.(2004)Ferns, Panangaden, and Precup] showed that bisimulation metrics can be computed as unique fixed points of a contraction map. A special case of this result for deterministic dynamics and on-policy samples was proved by [Castro(2020)]. Later, [Zhang et al.(2020)Zhang, McAllister, Calandra, Gal, and Levine] used bisimulation metrics to learn invariant observation embeddings for RL, where they proved these metrics can be computed while learning a policy. We will use these results in Section 4 to learn provably correct DFA embeddings. To our knowledge, no prior work considered using bisimulation metrics for learning task representations.
@@ 2 Background
Notation. Given a set $X$, we write $\mathcal{X}$ to denote a distribution over it, i.e., $\mathcal{X}\in\Delta(X)$, where $\Delta(X)\subset X\to[0,1]$ represents the set of all distributions over $X$. We use $\mathbb{I}\{\}$ for the event indicator function, where $\mathbb{I}\{p\}=1$ if and only if $p$ is true, and $\mathbb{I}\{p\}=0$ otherwise.
@@ 2.2 Deterministic Finite Automata
We use Deterministic Finite Automata (DFAs) with three-valued semantics to represent tasks.
Definition 2 (Deterministic Finite Automaton)
A Deterministic Finite Automaton (DFA) is defined as the tuple $\mathcal{A}=\langle Q,\Sigma,\delta,q_{0},F\rangle$, where $Q$ is the finite set of states, $\Sigma$ is the finite alphabet, $\delta:Q\times\Sigma\to Q$ is the transition function, where $\delta(q,a)=q^{\prime}$ denotes a transition to a state $q^{\prime}\in Q$ from a state $q\in Q$ by observing a symbol $a\in\Sigma$, $q_{0}\in Q$ is the initial state, and $F\subseteq Q$ is the set of final states. The three-valued semantics of a DFA is defined by a partition of its final states $F=F^{\top}\cup F^{\bot}$ and its extended (lifted) transition function $\delta^{*}:Q\times\Sigma^{*}\to Q$, where
@@
If $\delta^{*}(q_{0},w)\in F^{\top}$, then we say that $\mathcal{A}$ accepts $w$. Similarly, $\delta^{*}(q_{0},w)\in F^{\bot}$, then we say that $\mathcal{A}$ rejects $w$. $\mathcal{A}$ is called a plan DFA if its final states are sink states, i.e., $\forall q\in F,\forall a\in\Sigma,\delta(q,a)=q$.
DFAs can be minimized to a canonical form (up to a state isomorphism) using the algorithm of [Hopcroft(1971)], denoted by $\operatorname{minimize}(\mathcal{A})$. All minimized plan DFAs with nonempty accepted or rejected languages have single accepting or rejecting states, denoted by $\top$ or $\bot$, respectively, as their final states are all sink states. We denote the single-state accepting DFA by $\mathcal{A}_{\top}$ and the rejecting one by $\mathcal{A}_{\bot}$. Note that the approach presented in this paper is agnostic to the syntax of DFAs and can be trivially extended to other syntactic forms, e.g., compositional DFAs from [Yalcinkaya et al.(2024)Yalcinkaya, Lauffer, Vazquez-Chanlatte, and Seshia].
@@
In what follows, unless stated otherwise, all DFAs are plan DFAs.
We use DFAs to represent temporal tasks, which can be understood as plans. However, one can define multiple DFAs for the same task, i.e., DFAs without any additional assumptions are not canonical task representations. Therefore, we will need a notion of similarity between DFAs so that we can compare the tasks represented by them–we define bisimulation relation over DFAs next.
@@ 3 DFA-Conditioned Reinforcement Learning
We will define the DFA-conditioned RL problem over a distribution of DFAs, similar to the GCRL problem given in Appendix A. However, we need some extra structure over these DFAs since, in our case, goals are not simply sets of states but are DFAs encoding temporal tasks given to the policy.
@@
$\forall\mathcal{A}=\langle Q,\Sigma,\delta,q_{0},F\rangle\in D_{\Sigma,n},\forall w\in\Sigma^{*},\quad\operatorname{minimize}(\mathcal{A}^{\prime}=\langle Q,\Sigma,\delta,\delta^{*}(q_{0},w),F\rangle)\in D_{\Sigma,n}.$
@@
  • $D_{\Sigma,n}$, the set of DFAs, is the set of states,
  • $\Sigma$, the shared alphabet, is the set of actions,
@@
    $R_{D_{\Sigma,n}}(\mathcal{A}_{t},a)=\begin{cases}1&\text{if }T_{D_{\Sigma,n}}(\mathcal{A}_{t},a)=\mathcal{A}_{\top}\\ -1&\text{if }T_{D_{\Sigma,n}}(\mathcal{A}_{t},a)=\mathcal{A}_{\bot}\\ 0&\text{otherwise.}\end{cases}$
A DFA space is a set of DFAs closed under random walks, i.e., taking any random path in a DFA from this set and pruning the unreachable states results in a DFA in the set. In other words, one cannot get a DFA outside this set by taking a random walk with minimization, hence the name space.
@@
We now have all the theoretical machinery needed to formally state the DFA-conditioned RL problem. We first define the environment model and then continue with the statement of the problem.
@@
$\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}=\langle S\times D_{\Sigma,n},A,P_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}},R_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}},\iota_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}},\gamma\rangle$
@@
  • $S\times D_{\Sigma,n}$ is the (product) state space,
  • $A$ is the action space (of $\mathcal{M}$),
  • $P_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}:S\times D_{\Sigma,n}\times A\to\Delta(S\times D_{\Sigma,n})$ is the transition probability function defined by
    $P_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(s,\mathcal{A},a,s^{\prime},\mathcal{A}^{\prime})=P(s,a,s^{\prime})\mathbb{I}\left\{\mathcal{A}^{\prime}=T_{D_{\Sigma,n}}\left(\mathcal{A},L(s^{\prime})\right)\right\},$
  • $R_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}:S\times D_{\Sigma,n}\times A\to\{-1,0,1\}$ is the reward function defined by
    $R_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(s,\mathcal{A},a)=\begin{cases}1&\text{if }T_{D_{\Sigma,n}}(\mathcal{A},L(s^{\prime}))=\mathcal{A}_{\top}\\ -1&\text{if }T_{D_{\Sigma,n}}(\mathcal{A},L(s^{\prime}))=\mathcal{A}_{\bot}\\ 0&\text{otherwise,}\end{cases}$
@@
  • $\iota_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}\in\Delta(S\times D_{\Sigma,n})$ is the initial state distribution defined by
    $\iota_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(s,\mathcal{A})=\iota(s)\mathcal{D}_{\Sigma,n}(\mathcal{A}),$
@@
  • $\gamma\in[0,1)$ is the discount factor (of $\mathcal{M}$).
@@
Given an DFA-conditioned MDP $\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}$ as defined in 5, a DFA-conditioned policy is a mapping
@@
$J_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(\pi)=\mathbb{E}_{(s_{0},\mathcal{A}_{0})\sim\iota_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}}\left[\sum_{t=0}^{\mathcal{A}_{t}=\mathcal{A}_{\top}\lor\mathcal{A}_{t}=\mathcal{A}_{\bot}}\gamma^{t}\,R_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(s_{t},\mathcal{A}_{t},a_{t})\right],$
@@
$a_{t}\sim\pi(s_{t},\mathcal{A}_{t}),\quad s_{t+1}\sim P(s_{t},a_{t}),\quad\mathcal{A}_{t+1}=T_{D_{\Sigma,n}}\left(\mathcal{A}_{t},\,L(s_{t+1})\right),$
until the accepting or rejecting DFA is reached, i.e., $\mathcal{A}_{t}=\mathcal{A}_{\top}\lor\mathcal{A}_{t}=\mathcal{A}_{\bot}$. The objective is to solve
$\pi^{*}=\arg\max_{\pi}J_{\mathcal{M}\mid_{L}\mathcal{M}_{D_{\Sigma,n}}}(\pi),$
i.e., to learn a policy that maximizes the probability of satisfying a given temporal specification from $D_{\Sigma,n}$ (by driving its DFA representation to $\mathcal{A}_{\top}$) while operating in the underlying MDP $\mathcal{M}$.
Notice the difference between the GCRL problem formulation given in Appendix A and the DFA-conditioned one given above. Specifically, in GCRL, goals are static, i.e., they are given at the beginning, and the policy conditions on the same goal until it is accomplished. In the DFA-conditioned RL setting, through the labeling function, a given DFA task also evolves (based on the transition dynamics of its DFA space) as the policy interacts with the underlying MDP. Therefore, the policy is essentially learning to navigate two MDPs: the underlying MDP and the DFA space.
3.1 PAC-Learnability of the DFA-Conditioned Reinforcement Learning Problem
We show that the DFA-conditioned RL problem is Probably Approximately Correct (PAC)-learnable. To do so, we will use the PAC-MDP framework introduced by [Strehl et al.(2006)Strehl, Li, Wiewiora, Langford, and Littman]. An RL algorithm is called PAC-MDP (PAC in MDPs) if it finds a near-optimal policy with high probability in any MDP after a number of interactions that is polynomial in the problem’s key parameters, stated next.